Hershey is considering a new candy that will generate $30 million in sales and $20 million in cost of goods sold each year for the next 5 years (does not include depreciation expense). They will need a new machine that costs $20 million dollars. They will depreciate the machine to a zero book value using straight-line depreciation over 5 years. The machine can be sold at the end of the 5 years for $2,000,000. The tax rate is 40%. Determine the annual free cash flows.

Years 1 through 4: $3,600,000; year 5: $4,800,000

Years 1 through 4: $3,600,000; year 5: $5,600,000

Years 1 through 4: $7,600,000; year 5: $9,600,000

Years 1 through 4: $7,600,000; year 5: $8,800,000
Answer #1

Annual depreciation = 20,000,000 / 5 = 4,000,000

Years 1 through 4 = (Sales - costs - depreciation)(1 - tax) + depreciation

Years 1 through 4 = (30,000,000 - 20,000,000 - 4,000,000)(1 - 0.4) + 4,000,000

Years 1 through 4 = 3,600,000 + 4,000,000

Years 1 through 4 = $7,600,000

Year 5 cash flow = 7,600,000 + sale value - tax(sale value - book value)

Year 5 cash flow = 7,600,000 + 2,000,000 - 0.4(2,000,000 - 0)

Year 5 cash flow = 7,600,000 + 2,000,000 - 800,000

Year 5 cash flow = $8,800,000

Years 1 through 4: $7,600,000; year 5: $8,800,000